# Vision-Language-Action Models: Concepts, Progress, Applications and Challenges

### v1

@@ Abstract
Vision-Language-Action (VLA) models mark a transformative advancement in artificial intelligence, aiming to unify perception, natural language understanding, and embodied action within a single computational framework. This foundational review presents a comprehensive synthesis of recent advancements in Vision-Language-Action models, systematically organized across five thematic pillars that structure the landscape of this rapidly evolving field. We begin by establishing the conceptual foundations of VLA systems, tracing their evolution from cross-modal learning architectures to generalist agents that tightly integrate vision-language models (VLMs), action planners, and hierarchical controllers. Our methodology adopts a rigorous literature review framework, covering over 80 VLA models published in the past three years. Key progress areas include architectural innovations, parameter-efficient training strategies, and real-time inference accelerations. We explore diverse application domains such as humanoid robotics, autonomous vehicles, medical and industrial robotics, precision agriculture, and augmented reality navigation. The review further addresses major challenges across real-time control, multimodal action representation, system scalability, generalization to unseen tasks, and ethical deployment risks. Drawing from the state-of-the-art, we propose targeted solutions including agentic AI adaptation, cross-embodiment generalization, and unified neuro-symbolic planning. In our forward-looking discussion, we outline a future roadmap where VLA models, VLMs, and agentic AI converge to power socially aligned, adaptive, and general-purpose embodied agents. This work serves as a foundational reference for advancing intelligent, real-world robotics and artificial general intelligence.
@@
Vision-Language-Action , VLA , Artificial Intelligence , Robotics , Vision-Language Models , AI Agents , Agentic AI
journal: Proceedings of the IEEE
@@
Figure 1: Evolution from Isolated Modalities to Unified Vision-Language-Action Models. This figure illustrates the transition from separate vision, language, and action systems-each limited to its own domain-to integrated VLA models. VLA models enable robots to jointly perceive, understand language, and act, overcoming the fragmentation of earlier approaches and marking a major step toward adaptive, generalizable, and intelligent embodied agents.
@@ 1 Introduction
Before Vision-Language-Action (VLA) models were developed, progress in robotics and artificial intelligence happened mostly in separate areas: vision systems that could see and recognize images [44, 69], language systems that could understand and generate text [164, 137], and action systems that could control movement [49]. These systems worked well on their own, but they struggled to work together or handle new and unpredictable situations[46, 21]. As a result, ‘erstand complex environments or respond flexibly to real-world challenges.
@@
As illustrated in Figure 1, traditional computer vision models primarily based on convolutional neural networks (CNNs) , were tailored for narrowly specified tasks such as object detection or classification, requiring extensive labeled datasets and cumbersome retraining for even slight shifts in environment or objectives [156, 62]. These vision models could “see” (e.g., identifying apples in an orchard, as shown in Figure 1) but lacked any understanding of language or the ability to convert visual insights into purposeful actions. Language models, particularly large language models (LLMs), revolutionized text-based understanding and generation [23]; however, they remained restricted to processing language without the capability to perceive or reason about the physical world [76] (”Ripe apples in orchard” in Figure 1 exemplifies this limitation). Meanwhile, action-based systems in robotics, relying heavily on hand-crafted policies or reinforcement learning [122], enabled specific behaviors like object manipulation but demanded painstaking engineering and failed to generalize beyond narrowly scripted scenarios [119].
@@
Despite progress with VLMs, which achieved impressive multimodal understanding by combining vision and language [149, 25, 148], there remained a conspicuous integration gap: the inability to generate or execute coherent actions based on multimodal input [121, 107]. As further visualized in Figure 1 , most AI systems specialized at most in two modalities—vision-language, vision-action, or language-action—but struggled to fully integrate all three into a unified, end-to-end framework. Consequently, robots could recognize objects visually (”apple”), understand a corresponding textual instruction (”pick the apple”), or perform a predefined motor action (grasping), yet orchestrating these abilities into fluid, adaptable behavior was beyond reach. The result was a fragmented pipeline architecture that could not flexibly adapt to new tasks or environments, leading to brittle generalization and labor-intensive engineering efforts. This highlighted a critical bottleneck in embodied AI: without systems that could jointly perceive, understand, and act, intelligent autonomous behavior remained an elusive goal.
The pressing need to bridge these gaps catalyzed the emergence of VLA models. VLA models, conceptualized around 2021-2022, and pioneered by efforts such as Google DeepMind’s Robotic Transformer 2 (RT-2) [224], introduced a transformative architecture that unified perception, reasoning, and control within a single framework. As a solution to the limitations outlined in Figure 1, VLAs integrate vision inputs, language comprehension, and motor control capabilities, enabling embodied agents to perceive their surroundings, understand complex instructions, and execute appropriate actions dynamically. Early VLA approaches achieved this integration by extending vision-language models to include action tokens—numerical or symbolic representations of robot motor commands, thereby allowing the model to learn from paired vision, language, and trajectory data [121]. This methodological innovation dramatically improved robots’ ability to generalize to unseen objects, interpret novel language commands, and perform multi-step reasoning in unstructured environments [83].
VLA models represent a transformative step in the pursuit of unified multimodal intelligence, overcoming the long-standing limitations of treating vision, language, and action as separate domains [121]. By leveraging internet-scale datasets that integrate visual, linguistic, and behavioral information, VLAs empower robots to not only recognize and describe their environments but also to reason contextually and execute appropriate actions in complex, dynamic settings [196]. The progression illustrated in Figure 1 from isolated vision, language, and action systems to an integrated VLA paradigm-captures a fundamental shift toward the development of truly adaptive and generalizable embodied agents. Given the profound implications of this innovation, it is crucial to undertake a thorough and systematic review that draws from a comprehensive body of literature and critical analysis. First, such a review is necessary to clarify the foundational concepts and architectural principles that distinguish VLAs from their predecessors. Second, it provides a structured account of the rapid progress and key milestones in the field, enabling researchers and practitioners to appreciate the trajectory of technological advancements. Third, an in-depth review is essential for mapping the diverse range of real-world applications-from household robotics to industrial automation and assistive technologies-where VLAs are already demonstrating transformative potential. Furthermore, by critically examining the current challenges, such as data efficiency, safety, generalization, and ethical considerations, the review identifies barriers that must be addressed for widespread deployment. And Fifth, synthesizing these insights helps to inform the broader AI and robotics communities about emerging research directions and practical considerations, fostering collaboration and innovation.
@@
Concepts Foundation: What are VLAs Evolution Timeline Multimodal Integration Tokenization & Encoding Learning Paradigms Adaptive Execution
Figure 2: Mindmap for VLA Concepts. This diagram outlines the foundational components of Vision-Language-Action models, including their definitions, historical development, integration of multimodal signals, tokenization techniques, and adaptive execution. It sets the conceptual stage for understanding the structure and purpose of VLAs.
In this review, we systematically analyze the foundational principles, developmental progress, and technical challenges associated with VLA models. Our objective is to consolidate the current understanding of VLAs while identifying limitations and proposing future directions for their evolution. The review begins with a detailed examination of key conceptual foundations (Figure 2), including what constitutes a VLA model, its historical evolution, multimodal integration mechanisms, and language-based tokenization and encoding strategies. These conceptual components set the stage for understanding how VLAs are structured and function across modalities.
Building upon this, we present a unified view of recent progress and training efficiency strategies (Figure 3). This includes architectural innovations that have enabled more capable and generalizable VLA models, as well as data-efficient learning frameworks, parameter-efficient modeling techniques, and model acceleration strategies designed to reduce computational overhead without compromising performance. These advancements are critical for scaling VLA systems to real-world applications.
Following this, we delve into a comprehensive discussion of the current limitations faced by VLA systems (Figure 4). These include inference bottlenecks, safety concerns, high computational demands, limited generalization, and ethical implications. We not only highlight these pressing challenges but also provide an analytical discussion of potential solutions to address them.
Together, these three figures offer a visual framework that supports the textual analysis of this review. By outlining the conceptual landscape, recent innovations, and open challenges, this work aims to guide future research and encourage the development of more robust, efficient, and ethically grounded VLA systems.
Progress & Training Efficiency Architectural Innovations Applications of VLAs Data-Efficient Learning Parameter Efficiency Acceleration Methods
Figure 3: Mindmap – VLA Progress and Training Efficiency. This figure combines architectural advancements with training optimization methods such as data-efficient learning, parameter reduction, and model acceleration. It provides a visual summary of the technical progress driving the scalability and real-world deployment of VLAs.
@@
Challenges Inference Constraints Multimodal Action Safety & Bias System Complexity Compute Demands Generalization Gaps Environment Robustness Ethics & Society
Figure 4: Mindmap – VLA Challenges. This diagram highlights key barriers to robust VLA deployment, including inference limitations, bias, system complexity, generalization gaps, and ethical concerns. It also motivates the need for innovative solutions and future research directions to overcome these challenges.
@@ 2 Concepts of Vision-Language-Action Models
VLA models represent a new class of intelligent systems that jointly process visual inputs, interpret natural language, and generate executable actions in dynamic environments. Technically, VLAs combine vision encoders (e.g., CNNs, ViTs), language models (e.g., LLMs, transformers), and policy modules or planners to achieve task-conditioned control. These models typically employ multimodal fusion techniques—such as cross-attention, concatenated embeddings, or token unification—to align sensory observations with textual instructions. Unlike traditional visuomotor pipelines, VLAs support semantic grounding, enabling context-aware reasoning, affordance detection, and temporal planning. A typical VLA model observes the environment through camera or sensor data, interprets goals expressed in language (e.g., “pick up the red apple”) (Figure 5), and outputs low-level or high-level action sequences. Recent advancements integrate imitation learning, reinforcement learning, or retrieval-augmented modules to improve sample efficiency and generalization. This review examines how VLA models have evolved from foundational fusion architectures to general-purpose agents capable of real-world deployment across robotics, navigation, and human-AI collaboration.
VLA models are multimodal artificial intelligence systems that unify visual perception, language comprehension, and physical action generation into a single framework. These models enable robots or AI agents to interpret sensory inputs (e.g., images, text), understand contextual meaning, and autonomously execute tasks in real-world environments-all through end-to-end learning rather than isolated subsystems. As shown conceptually in Figure 5, VLA models bridge the historical disconnect between visual recognition, language comprehension, and motor execution that limited the capabilities of earlier robotic and AI systems.
@@ 2.1 Evolution and Timeline
    Foundational Integration (2022–2023). Early VLAs established basic visuomotor coordination through multimodal fusion architectures. [157] first combined CLIP embeddings with motion primitives, while [141] demonstrated generalist capabilities across 604 tasks. [18] achieved 97% success rates in manipulation through scaled imitation learning, and [86] introduced temporal reasoning via transformer-based planners. By 2023, [224] enabled visual chain-of-thought reasoning, and [34] advanced stochastic action prediction through diffusion processes. These foundations addressed low-level control but lacked compositional reasoning [216], prompting innovations in affordance grounding [78].
@@
    Specialization and Embodied Reasoning (2024). Second-generation VLAs incorporated domain-specific inductive biases. [202] enhanced few-shot adaptation through retrieval-augmented training, while [210] optimized navigation via 3D scene-graph integration. [39] introduced reversible architectures for memory efficiency, and [183] addressed partial observability with physics-informed attention. Simultaneously, [5] improved compositional understanding through object-centric disentanglement, and [220] extended applications to autonomous driving via multi-modal sensor fusion. These advances required new benchmarking methodologies [196].
@@
    Generalization and Safety-Critical Deployment (2025). Current systems prioritize robustness and human alignment. [205] integrated formal verification for risk-aware decisions, while [42] demonstrated whole-body control through hierarchical VLAs. [19] optimized compute efficiency for embedded deployment, and [102] combined neural-symbolic reasoning for causal inference. Emerging paradigms like [100]’s affordance chaining and [13]’s sim-to-real transfer learning address cross-embodiment challenges, while [108] bridges VLAs with human-in-the-loop interfaces through natural language grounding.
Figure 6 presents a comprehensive timeline highlighting the evolution of 47 VLA models developed between 2022 and 2025. The earliest VLA systems, including CLIPort [157], Gato [141], RT-1 [18], and VIMA [86], laid the foundation by combining pretrained vision-language representations with task-conditioned policies for manipulation and control. These were followed by ACT [216], RT-2 [224], and VoxPoser [78], which integrated visual chain-of-thought reasoning and affordance grounding. Models like Diffusion Policy [34] and Octo [167] introduced stochastic modeling and scalable data pipelines. In 2024, systems such as Deer-VLA [202], ReVLA [39], and Uni-NaVid [210] added domain specialization and memory-efficient designs, while Occllama [183] and ShowUI [108] tackled partial observability and user interaction. The trajectory continued with robotics-focused VLAs like Quar-VLA [43] and RoboMamba [111]. Recent innovations emphasize generalization and deployment: SafeVLA [205], Humanoid-VLA [42], and MoManipVLA [190] incorporate verification, full-body control, and memory systems. Models such as Gr00t N1 [13] and SpatialVLA [136] further bridge sim-to-real transfer and spatial grounding. This timeline illustrates how VLAs have advanced from modular learning to general-purpose, safe, and embodied intelligence.
20222023 CLIPort [157] Gato [141] RT-1 [18] VIMA [86] ACT [216] RT-2 [224] VoxPoser [78] Diffusion Policy [34] Octo [167] OpenVLA [94] 20242024 Deer-VLA [202] Uni-NaVid [210] ReVLA [39] Occllama [183] Pi-0 [14] RDT-1B [112] CogAct [102] EdgeVLA [19] ShowUI [108] NaviLa [32] Quar-VLA [43] Bi-VLA [59] RoboMamba [111] Otter [75] PointVLA [96] CombatVLA [29] HybridVLA [110] 20252025 CoVLA [5] OpenDriveVLA [220] ORION [56] ObjectVLA [223] ConRFT [31] Hi Robot [155] TLA [70] RaceVLA [153] DexVLA [185] Humanoid-VLA [42] SafeVLA [205] MoManipVLA [190] VLA-Cache [195] TinyVLA [186] Gr00t N1 [13] NORA [79] SpatialVLA [136] MoLe-VLA [213]
Figure 6: Comprehensive timeline of Vision-Language-Action models (2022–2025), showing evolution from foundation to 45 specialized VLA systems. Organized chronologically with thematic grouping.

### v2

# Vision-Language-Action (VLA) Models: Concepts, Progress, Applications and Challenges

@@ Abstract
Vision-Language-Action (VLA) models mark a transformative advancement in artificial intelligence, aiming to unify perception, natural language understanding, and embodied action within a single computational framework. This foundational review presents a comprehensive synthesis of recent advancements in Vision-Language-Action models, systematically organized across five thematic pillars that structure the landscape of this rapidly evolving field. We begin by establishing the conceptual foundations of VLA systems, tracing their evolution from cross-modal learning architectures to generalist agents that tightly integrate vision-language models (VLMs), action planners, and hierarchical controllers. Our methodology adopts a rigorous literature review framework, covering over 80 VLA models published in the past three years. Key progress areas include architectural innovations, efficient training strategies, and real-time inference accelerations. We explore diverse application domains such as autonomous vehicles, medical and industrial robotics, precision agriculture, humanoid robotics, and augmented reality. The review further addresses major challenges across real-time control, multimodal action representation, system scalability, generalization to unseen tasks, and ethical deployment risks. Drawing from the state-of-the-art, we propose targeted solutions including agentic AI adaptation, cross-embodiment generalization, and unified neuro-symbolic planning. We outline a forward-looking roadmap where VLA models, VLMs, and agentic AI converge to strengthen socially aligned, adaptive, and general-purpose embodied agents. This work, therefore, is expected to serve as a foundational reference for advancing intelligent, real-world robotics and artificial general intelligence. The project repository is available on GitHub(Source Link)
@@
Vision-Language-Action , Action Tokenization , Artificial Intelligence , Robotics , Vision-Language Models
journal: Information Fusion
@@ 1 Introduction
Before Vision-Language-Action (VLA) models were developed, progress in robotics and artificial intelligence happened mostly in separate domains: vision systems that could acquire, interpret and recognize images [55, 87, 176], language systems that could understand and generate text [213, 177], and action systems that could control movement [60]. These isolated systems worked well on their own, but struggled to work together and generalize to novel scenarios or adapt to the complexity and unpredictability of real-world challenges [57, 22].
Figure 1: Evolution from isolated modalities to unified Vision–Language–Action models. Integrated perception, language, and action enable adaptive, generalizable embodied intelligence.
@@
As illustrated in Figure 1, traditional computer vision models, primarily based on convolutional neural networks (CNNs), were tailored for narrowly specified tasks such as object detection [25, 157, 26, 24, 174] or classification [205, 89, 95, 73], requiring extensive labeled datasets and cumbersome retraining for even slight shifts in environment or objectives [200, 77]. These vision models could “see” (e.g., identifying apples in an orchard, as shown in Figure 1) but lacked any understanding of language or the ability to convert visual insights into desired actions. Language models, particularly large language models (LLMs), revolutionized text-based understanding and generation [28]; however, they remained restricted to processing language without the capability to perceive or reason about the physical world [98] (“Ripe apples in orchard” in Figure 1 exemplifies this limitation). Meanwhile, action-based systems in robotics, relying heavily on hand-crafted policies or reinforcement learning [158], enabled specific behaviors like object manipulation but demanded painstaking engineering and failed to generalize beyond specifically designed scenarios [153].
@@
Despite progress with VLMs, which achieved impressive multi-modal understanding by combining vision and language [190, 30, 298, 37, 10, 189], there remained a clear integration gap: the inability to generate or execute coherent actions based on multi-modal input [156, 137]. As further visualized in Figure 1 , most AI systems are specialized in one or two modalities such as vision-language, vision-action, or language-action, which struggled to fully integrate all three into a unified, end-to-end frameworks. Consequently, robots could recognize objects visually (“apple”), understand a corresponding textual instruction (“pick the apple”), or perform a predefined motor action (grasping), yet integrating and performing all these abilities into fluid, adaptable behavior has been missing. The result was a pipeline that could not flexibly adapt to new tasks or environments, leading to brittle generalization and labor-intensive engineering efforts. This limitation highlighted a critical bottleneck in embodied AI: without systems that could jointly perceive, understand, and act, intelligent autonomous behavior remained a challenging goal.
@@
VLA Concepts Foundation: What are VLAs? Evolution Timeline Multimodal Integration Tokenization & Encoding Learning Paradigms Adaptive Execution
Figure 2: Mind map of core VLA concepts. Each color-coded branch highlights a foundational dimension: definitions (foundation), historical evolution, multimodal integration, tokenization and encoding, learning paradigms, and adaptive execution in embodied settings.
@@
Vision-Language-Action (VLA) Models VLA Progress & Training Efficiency Architectural Innovations VLA Applications Data-Efficient Learning Parameter Efficiency Acceleration Methods VLA Challenges Inference Constraints Multimodal Action Safety & Bias System Complexity Compute Demands Generalization Gaps Environmental Robustness Ethics & Society
Figure 3: Mind map illustrating VLA model ecosystem: progress in training efficiency (architectural innovations, data/parameter efficiency, acceleration) alongside key challenges (inference constraints, multimodal action, safety, generalization, ethics) that must be overcome for scalable real-world deployment.
The pressing need to bridge these gaps catalyzed the emergence of VLA models. VLA models, conceptualized around 2021-2022, and pioneered by efforts such as Google DeepMind’s Robotic Transformer 2 (RT-2) [299], introduced a transformative architecture that unified perception, reasoning, and control within a single framework. As a solution to the limitations outlined in Figure 2, VLAs integrate vision inputs, language comprehension, and motor control capabilities, enabling embodied agents to perceive their surroundings, understand complex instructions, and execute appropriate actions dynamically. Early VLA approaches achieved this integration by extending vision-language models to include action tokens numerical or symbolic representations of robot motor commands, thereby allowing the model to learn from paired vision, language, and trajectory data [156]. This methodological innovation dramatically improved robots’ ability to generalize to unseen objects, interpret novel language commands, and perform multi-step reasoning in unstructured environments [107].
VLA models represent a transformative step in the development of unified multi-modal intelligence, overcoming the long-standing limitations of treating vision, language, and action as separate domains [156]. By leveraging internet-scale datasets that integrate visual, linguistic, and behavioral information, VLAs empower robots to not only recognize and describe their environments but also to reason contextually and execute appropriate actions in complex, dynamic settings [254]. The progression illustrated in Figures 2 and 3 from isolated vision, language, and action systems to an integrated VLA paradigm-captures a fundamental shift toward the development of truly adaptive and generalizable embodied agents. In light of the transformative potential of this paradigm, a comprehensive and critically informed review of the current literature is both timely and essential. First, such a review is necessary to clarify the foundational concepts and architectural principles that distinguish VLAs from their predecessors. Second, it provides a structured account of the rapid progress and key milestones in the field, enabling researchers and practitioners to appreciate the trajectory of algorithmic and technological advancements. Third, an in-depth review is essential for mapping the diverse range of real-world applications - from household robotics to industrial automation and assistive technologies-where VLAs are already demonstrating transformative potential. Furthermore, by critically examining the current challenges, such as data efficiency, safety, generalization, and ethical considerations, the review identifies barriers that must be addressed for widespread deployment. Finally, synthesizing these insights helps to inform the broader AI and robotics communities about emerging research directions and practical considerations, fostering collaboration and innovation.
In this review, we systematically analyze the foundational principles of VLA models. Additionally, we discuss their developmental progress and technical challenges. Our objective is to consolidate the current understanding and applications of VLAs while identifying limitations and proposing future directions for their evolution.The review begins with a detailed examination of key conceptual foundations (Figure 2), including the definition of VLA models, their historical evolution, mechanisms for multimodal integration, and unified tokenization and representation strategies spanning vision, language, and action. These conceptual descriptions set the stage for understanding how VLAs are structured and function across modalities.
Building upon this description, we present a unified view of recent progress and training efficiency strategies (Figure 3). This includes architectural innovations adopted and extended within VLA models, along with data-efficient learning frameworks, parameter-efficient modeling techniques, and model acceleration strategies originally developed in broader machine learning and robotics contexts. Together, these advances are critical for scaling VLA systems to real-world applications.
Following this, we present a comprehensive discussion of the current limitations encountered by VLA systems (Figure 3), many of which reflect broader challenges in embodied AI and robotics but arise in distinct and compounded forms due to the tight integration of vision, language, and action. The limitations to be discussed include inference bottlenecks, safety concerns, high computational demands, limited generalization, and ethical implications. We not only highlight these pressing challenges but also provide an analytical discussion on potential solutions to address them.
Together, these three figures offer a visual depiction that lays out the framework and supports the textual analysis presented in this review. By outlining the conceptual landscape, recent innovations, and open challenges, this work aims to guide future research and encourage the development of more robust, efficient, and ethically grounded VLA systems.
Figure 4 summarizes the overall structure and logical flow of this review and illustrates how the manuscript is organized to provide a comprehensive and systematic analysis of VLA research. As depicted in the figure, the paper progresses from foundational concepts to recent advances, applications, challenges, and future research directions, ensuring a coherent narrative across sections. To construct this architecture, an extensive and rigorous literature search was conducted using the primary keywords “Vision–Language–Action” and “Vision–Language Models,” together with the commonly used abbreviation “VLA.” These keywords were employed to retrieve candidate studies from major academic and technical repositories, including Hugging Face, arXiv, ScienceDirect, Nature, IEEE Xplore, Wiley, and Springer Nature. The resulting corpus was further refined through manual screening to ensure relevance, technical depth, and alignment with the scope of this review. Only articles that directly contributed to the understanding of VLA concepts, methodological progress, application domains, and open challenges were retained. This multi-stage filtering process enabled a balanced coverage of both foundational and state-of-the-art works while avoiding peripheral or loosely related studies. As a result, Figure 4 not only reflects the thematic organization of the paper but also embodies the underlying review methodology.
This paper follows a structured, hierarchical organization that systematically develops the foundations, evolution, and implications of VLA models, as summarized in Figure 4. The presentation begins with an introduction that motivates embodied intelligence and the emergence of VLAs, followed by a concepts section that establishes core principles including multimodal integration, tokenization, learning paradigms, and real-time control. Building on these foundations, the progress section examines architectural innovations, training and efficiency advancements, and parameter-efficient acceleration strategies. The applications section then grounds these developments in real-world domains, including humanoid robotics, autonomous vehicles, healthcare, agriculture, industrial systems, and interactive AR navigation. This is followed by a focused analysis of key challenges, including real-time inference, safety, generalization, system integration, and ethical considerations. The paper concludes with a future roadmap that distills cross-cutting research directions in continual learning, scalability, interpretability, and embodied intelligence.
{forest}
Figure 4: Flow and structure of this paper, from conclusion and future roadmap back through applications, progress, concepts, to introduction.
@@ 2 Concepts of Vision-Language-Action Models
VLA models represent a class of intelligent systems that jointly process visual inputs, interpret natural language instructions, and generate executable action representations that can be instantiated on physical robotic hardware operating in dynamic environments. Technically, VLAs combine vision encoders (e.g., CNNs, ViTs), language models (e.g., LLMs, transformers), and policy modules or planners to achieve task-conditioned control. These models typically build upon multimodal fusion techniques established in vision-language models, such as cross-attention, concatenated embeddings, or token unification, and extend them to align sensory observations, linguistic instructions, and action representations.
Unlike traditional visuomotor pipelines, VLAs support semantic grounding [154], enabling context-aware reasoning [228], affordance detection [79], and temporal planning [141]. A typical VLA model observes the environment through camera or sensor data, interprets goals expressed in language (e.g., “pick up the red apple”) (Figure 5), and outputs low-level or high-level action sequences that could be implemented by automated systems to perform the action. Recent advancements integrate imitation learning, reinforcement learning, or retrieval-augmented modules to improve sample efficiency and generalization. This review examines how VLA models have evolved from foundational fusion architectures to general-purpose agents capable of real-world deployment across robotics, navigation, and human-AI collaboration.
VLA models are multi-modal artificial intelligence systems that unify visual perception, language comprehension, and physical action generation into a single framework. These models enable robots or AI agents to interpret sensory inputs (e.g., images, text), understand contextual meaning, and autonomously execute tasks in real-world environments - all through end-to-end learning and action rather than isolated subsystems. As shown conceptually in Figure 5, VLA models bridge the historical disconnect between visual recognition, language comprehension, and/or motor execution that limited the capabilities of earlier robotic and AI systems.
@@ 2.1 Evolution and Timeline
    Foundational Integration (2022–2023). Early VLAs established basic visuomotor coordination through multi-modal fusion architectures. [202] first combined CLIP embeddings with motion primitives, while [181] demonstrated generalist capabilities across 604 tasks. [19] achieved 97% success rates in manipulation through scaled imitation learning, and [112] introduced temporal reasoning via transformer-based planners. By 2023, [299] enabled visual chain-of-thought reasoning, and [40] advanced stochastic action prediction through diffusion processes. These foundations addressed low-level control but lacked compositional reasoning, the ability to decompose complex tasks into reusable, semantically grounded sub-actions and recombine them across novel contexts-prompting subsequent innovations in affordance grounding [287, 100].
@@
    Specialization and Embodied Reasoning (2024). Second-generation VLAs incorporated domain-specific inductive biases. [269] enhanced few-shot adaptation through retrieval-augmented training, while [280] optimized navigation via 3D scene-graph integration. [48] introduced reversible architectures for memory efficiency, and [239] addressed partial observability with physics-informed attention. Simultaneously, [5] improved compositional understanding through object-centric disentanglement, and [293] extended applications to autonomous driving via multi-modal sensor fusion. However, these advances required new benchmarking methodologies [254].
@@
    Generalization and Safety-Critical Deployment (2025). Latest systems prioritize robustness and human alignment. [274] integrated formal verification for risk-aware decisions, while [53] demonstrated whole-body control through hierarchical VLAs. [20] optimized compute efficiency for embedded deployment, and [131] combined neural-symbolic reasoning for causal inference. Emerging paradigms like [129]’s affordance chaining and [14]’s sim-to-real transfer learning address cross-embodiment challenges, while [139] bridges VLAs with human-in-the-loop interfaces through natural language grounding.
Figure 6 presents a comprehensive timeline highlighting the evolution of 45 VLA models developed between 2022 and 2025. The earliest VLA systems, including CLIPort [202], Gato [181], RT-1 [19], and VIMA [112], laid the foundation by combining pretrained vision-language representations with task-conditioned policies for manipulation and control. These early VLA systems were followed by ACT [287], RT-2 [299], and VoxPoser [100], which integrated visual chain-of-thought reasoning and affordance grounding. Models like Diffusion Policy [40] and Octo [218] introduced stochastic modeling and scalable data pipelines. In 2024, systems such as Deer-VLA [269], ReVLA [48], and Uni-NaVid [280] added domain specialization and memory-efficient designs, while Occllama [239] and ShowUI [139] tackled partial observability and user interaction. The trajectory continued with robotics-focused VLAs like Quar-VLA [54] and RoboMamba [143]. Recent innovations emphasize generalization and deployment: SafeVLA [274], Humanoid-VLA [53], and MoManipVLA [246] incorporate verification, full-body control, and memory systems. Models such as Gr00t N1 [14] and SpatialVLA [175] further bridge sim-to-real transfer and spatial grounding. This timeline illustrates how VLAs have advanced from modular learning to general-purpose, safe, and embodied intelligence.
2022 (Foundation)2023 (Scaling / Policies)2024 (Specialization)2025 (Generalization / Advanced) CLIPort [202] Gato [181] RT-1 [19] VIMA [112] ACT [287] RT-2 [299] VoxPoser [100] Diffusion Policy [40] Octo [218] OpenVLA [122] Deer-VLA [269] Uni-NaVid [280] ReVLA [48] Occllama [239] Pi-0 [15] RDT-1B [145] CogAct [131] EdgeVLA [20] ShowUI [139] NaviLa [38] Quar-VLA [54] Bi-VLA [74] RoboMamba [143] Otter [97] PointVLA [125] HybridVLA [142] CoVLA [5] OpenDriveVLA [293] ORION [69] ObjectVLA [297] ConRFT [36] Hi Robot [199] TLA [88] RaceVLA [195] DexVLA [241] Humanoid-VLA [53] SafeVLA [274] MoManipVLA [246] VLA-Cache [252] TinyVLA [242] Gr00t N1 [14] NORA [103] SpatialVLA [175] MoLe-VLA [283] Long-VLA [64] RetoVLA [123] Vlaser [256] Discrete Diffusion VLA [138] Being-H0 [152] EgoVLA [257] StereoVLA [47] GeoVLA [212] EfficientVLA [260]
Figure 6: Comprehensive timeline of Vision–Language–Action (VLA) models (2022–2025), organized by year with high-contrast color coding and thematic grouping.